what is the length of the longest stick which can fit in the bottom of a rectangular box if its base is 25 inches long and 10 inches wide
Answer by Cromlix(4381)

Hi there,
With a rectangular box measuring 25 inches
by 10 inches. A stick could lie along the diagonal
of the box.
The diagonal^2 = Length^2 + width^2
Diagonal squared = 25^2 + 10^2
Diagonal squared = 625 + 100
Diagonal = square root of 725
Diagonal = 26.9 (1 decimal place) inches.
This is the size of the biggest stick which
can fit in the box.
Hope this helps. :-)